An object moving with a speed of υ, = 8 m/s and accelerates due to a force exerted on it. After 5 s the speed of the object is 14 m/s. Calculate the displacement of the object.
1
Expert's answer
2020-03-19T11:30:34-0400

For motion with constant acceleration:


"d=\frac{v_i+v_f}{2}t"

"d=\frac{8+14}{2}(5)=55\ m"